1. Solve the nonhomogeneous linear system of differential equations d / 8 3 -10) dt ( 3 8 94 25 u = Ju+ 10 ) 1 -25
Answer #1

non-homagencous uncor system of differential quaton is given by Th a 5 16 3 3 Here u is such that a 2X1 vector The be rewTite

Cven by the charat teristte equatlon is m-16m+55 = 0 -0 m-1n- 5 m + 55 m(m-1) 5m -1) = O (m-1) m-5) m m-5 Thr compumentary so

Diffeventiqting we have cuith espeot to t dui dt A et58 e5t- 5 .(vi) From ) have, 342du dt 10 25 8uI 3u2 A e5Best.5 -8 aeua